### v1

@@ 2.5 Message-passing Neural Network
where $\{x_{i},x_{j}\}\in A=1$ and $msg(\cdot)$ is a message function with the source and target nodes as inputs. From Eq.3, the node feature of $x_{i}$ is updated with the computed message $(x_{i},x_{j})$, and together linearly transformed into latent space, producing the latent feature $h_{i}\in\mathcal{H}$.
Eq. 4 present the message-passing neural network in vector view as follow:
@@
A global readout function (i.e., mean or max pooling) summarizes the latent node-wise features ($\mathcal{H}$) and computes a global graph embedding for the input graph by non-linearly transforming the pooled feature.
Hence, we propose a Messpage-passing neural network for Polygon geometries, PolyMP, as shown in Fig. 1. The PolyMP considers the point coordinates of polygons as input node features and the exterior connectivity of polygons as input edges. The message-passing layers of PolyMP then compute the “message” for each node according to the graph adjacency. Following existing works on point clouds [36, 11] and polygons Liu et al. [35], we compute the message of relative positions of two connected nodes: $msg(x_{i},x_{j})=|x_{j}-x_{i}|$, which enable aggregating the local geometric features of polygons for updating node features. After message-passing layers, a global pooling layer is applied over the polygons with aggregated node features to generate a whole graph embedding, which is then used for the downstream shape classification of polygons.
@@
Figure 1: Model architecture of PolyMP. PolyMP consists of two message-passing layer and a classification layer (Multi-layer perceptron).
3 Experiment
3.1 Datasets
3.1.1 Glyph Dataset
We introduce a synthetic dataset of highly variable geometric shapes for benchmarking the classification performance of learning models on vector polygons. The dataset consists of 26 Latin alphabet character glyph geometries (here also equal to semantic classes A to Z) of fonts gathered from an online source [40]. An equivalent approach to construct a rich dataset of geometric shapes has been used to test algorithms in computational geometry before [41].
We extract the boundaries of glyphs ( i.e., contour lines, as called in typography) for 1,413 sans serif and 1,002 serif fonts, to produce 2D simple polygon geometries compliant with OpenGeospatialConsortium [38]. Serif and sans serif fonts are the two main typographic font families. Serif glyphs have salient decorative strokes ending the main character strokes increasing legibility of body text, whereas sans serif glyphs have clean strokes suitable e.g., for the display of headers. Importantly, these minor variations do not alter the Gestalt of the shapes (as proven by reader’s ability to consistently read the symbols). Thus, the labels assigned to the symbols are stable and robust. We encode each polygon geometry into a fixed-size feature matrix $\in R^{n\times 3}$ where each feature vector $\in R^{3}$ includes the 2D coordinates of vertices $(x,y)$ as geometric feature and a binary feature $(0,1)$ to indicate the positions of 2D coordinates, whether on outer rings or holes of polygons. Examples of glyph geometry are displayed in Figure 2.
3.1.2 OSM Dataset
Latin alphabet glyph geometries share geometric similarity with building footprints. We use a building geometry dataset designed to test the generalizability of learning models on spatial objects proposed by Yan et al. [6]. The dataset contains 10,000 real-world building footprints extracted from OSM [23], labelled by 10 categories based on a template matching to letters [42]. The building footprints are randomly rotated and reflected from the original canonical samples (Figure 3). This is currently the most comprehensive real-world dataset of vector geometries with purely shape-based labels (as opposed to, e.g., building use, which should not be used for shape classification). We note that the assignment of labels in this dataset is somewhat subjective and likely less robust than in the glyph-based dataset. Note, in particular, that shape reflection is problematic for label stability, e.g., for asymmetric letters ($R$), or mirror shapes of distinct labels ($S$,$Z$).
3.1.3 Pre-processing
Label-preserving transformations
To examine the invariances of geometric transformations of polygon representations and corresponding learning models with respects to data transformations, we consider four training datasets under data transformations, each consisting of 20%, 40%, 60% and 80% of glyph polygons that have undergone label-preserving geometric transformations. The details of transformations are presented in Table 1. The original and transformed data samples are visualised in Fig. 4. We thus obtain five training datasets, including the original dataset: dataset 0%, 20%, 40%, 60% and 80% for the experiment. The test set consists of data randomly sampled from the five datasets.
Transformation Operation
Polygon Simplifications
To test the invariances to trivial vertices of polygon, we apply the Douglas–Peucker algorithm [43] (tolerance = 1.0) on the original size of glyph polygon (in the order of 50 $\times$ 50 units) to generate simplified samples for glyph dataset. The resulting polygon linestrings are defined by fewer vertices (in particular, excluding all co-linear or close to co-linear vertices). Notice that the simplification occurs before the normalisation to the range of (-1,1) and any augmentation of polygons, ensuring that the simplified geometries preserve the topology and essence of the shape of the original glyph (see Fig. 5).
3.2 Training Configuration
We evaluate the performances of learning models from previous studies, in particular: the VeerCNN in Veer et al. [2], based on 1D sequences converted from polygons as inputs; the DeepSet from Zaheer et al. [9] and SetTransformer Lee et al. [12], based on polygon vertices as input point clouds; and the GCAE in Yan et al. [6], representing polygons as input graphs. We then compare their performances with the proposed message-passing model for polygons (PolyMP).
The model configurations used in our experiments are in Table 2. Correspondingly, we depict the data representations of polygons and baseline models in Figure 6. The proposed PolyMP is demonstrated in Figure 1. For experimental control, every model tested was constrained to a comparable model depth and latent space of feature embedding.
For model training on Glyph dataset and fine-tuning on OSM dataset, we compute the cross entropy loss [44] to determine training gradients for fast training convergence and good model generalizability. We use the Adam optimization [45] with an initial learning rate of $0.01$. For better model generalization, we adopt a learning rate decay strategy that reduces the learning rate by a factor of $10$ once the model training reaches a performance plateau over $25$ epochs. An early-stopping mechanism halts the training once the loss ceases to decrease for $50$ epochs, to mitigate overfitting. We set the batch size to $64$ and the total training epochs to $100$. Thus, importantly, the models here are not designed to individually outperform current state-of-the-art for each architecture, but to maintain comparability across architectures and to enable evaluation of the effects of data representations on the results.
@@ 3.2 Training Configuration
PolyMP (Ours) 2 {2, 64, 64} 11,837 Global avg.
@@ 4.1 Results on Glyph Dataset
Table 3 and 4 demonstrates the model performances testing on Glyph dataset and the simplified polygons, respectively.
At $0\%$ transformation ratio in Table 3, we observe that the PolyMP achieves $99.41\%$, $42.91\%$ and $80.03\%$ acc. on the original, rotated and scaled test samples, whereas VeerCNN achieves $59.07\%$ acc. on sheared samples. While deepset, set transformer and GCAE maintain comparable performances ($94.62\%$, $65.01\%$; and $96.25\%$, $67.91\%$; and $95.16\%$, $64.54\%$, respectively) on original and scaled test samples, the models suffer large performance deterioration, achieving only $23.67\%$, $26.00\%$ and $25.23\%$, $27.79\%$ and $31.38\%$ and $38.84\%$ test acc. on rotated and sheared samples, respectively. From the overall accuracy, the VeerCNN outperforms the PolyPM by a slight margin ($69.43\%$ and $69.01\%$).
At 20% transformation ratio, the PolyMP improves model test acc. on rotated, scaled and sheared samples rapidly, achieving 83.17% (+40.26%), 91.29% (+11.26%) and 86.86% (+33.54%) respectively. In overall acc., the PolyMP achieves 90.12% O.A. (+21.11%) and outperforms VeerCNN. In comparison, the other models can not match with PolyMP and achieve comparable performances from increasing data diversity of training set.
At 40% and 60% transformation ratio, we notice that the accuracy improvements gained from increasing the ratio of geometric transformed samples in training set have converged. In this case, the PolyMP maintains its leading performances among models (91.91% and 93.06% O.A.), and VeerCNN (83.30% and 84.88% O.A.), deepset (80.12% and 81.49% O.A.), set transformer (80.33% and 81.94% O.A.) and GCAE (72.72% and 75.50% O.A.) fall behind.
At 80% transformation ratio, the PolyMP achieves the optimal test performances at rotated and sheared samples, giving 90.35% and 91.52% test accuracy. This empirical results suggest that compared to the sequence encoding of polygons and CNN as well as the set representation and set-based learning models, the graph representation together with message-passing neural network (i.e., PolyMP) are robust to the permutations of feed-in order of polygon vertices, which occurs in geometric transmutations of polygons, in particular rotation and shearing.
From a different perspective, the ratio of original samples is decreasing in the training set progressively from the 0% to 80% transformation ratios datasets. The PolyMP only encounters 1.46% test performance deterioration along this progression (99.41% to 97.95% acc.), while the VeerCNN (-4.13% acc.), deepset (-3.96% acc.), set transformer (-5.08% acc.) and GCAE (-5.58% acc.) suffer larger performance drops on the test accuracy of original samples.
@@
PolyMP 97.95 90.35 93.30 91.52 93.42
In Table 4, the test performances of models trained and evaluated on the simplified polygons of Glyph dataset are presented. In this experiment setting, we focus on models’ test performances on simplified polygons containing fewer numbers of vertices, and compare them with the test performances showed in Table 3.
We observe similar test performance trends to those of models evaluated on the simplified polygons. Noticeably, there is a significant performance deterioration in VeerCNN compared with the test performances in Table 3 on simplified polygons. Specifically, at 0% transformation ratio, the test performances of VeerCNN drops from 69.43% O.A. (Table 3) to 66.56% O.A. (Table 4), i.e., a 2.87% performance deterioration. As the data diversity increases, the performance deterioration in VeerCNN increases, too, from 5.2%, progressively through to 5.56% O.A. deterioration respectively from 20% to 80% transformation ratios.
The performance deterioration of VeerCNN strongly indicates that CNN models are not robust to varying the numbers of vertices in polygons. This can be attributed to a number of factors, incl. the data encoding of polygons in VeerCNN as as fixed-length 1D sequences with zero padding. Simplifying polygons and removing trivial vertices from polygons alters the length of 1D sequences, and hence the geometric information of simplified polygons are not only diluted by the zero padding in the 1D sequences, but also by alteration of vertex adjacency explicitly encoded in the 1D sequence.
In comparison, we notice only small performance drops occur in general set-based and graph-based learning models. We credit this to set and graph representation of polygons and the corresponding learning models as both set and graph-based models accept input polygons of varying lengths. In particular, the graph-based message-passing network (i.e., PolyMP) learns salient local node features from connected neighbours in the message-passing function with permutation-invariant pooling operations (i.e., Max-pooling), as shown in E.q. 4. A similar pooling operation is also adopted in set-based learning models [9, 12].
@@
PolyMP 97.01 90.19 91.53 90.31 92.22
@@ 4.2 Results on OSM Dataset
We fine-tune models pre-trained on the glyph 80% dataset with the OSM dataset, and report the classification test accuracy of fine-tuned models in Table 5 for building footprints with original and normalised coordinates respectively.
PolyMP achieves optimal test performance on normalised footprints with $91.92\%$ O.A. on the OSM dataset, and performs better than all other methods. Both set-based methods (i.e., DeepSet: $95.71\%$ and SetTransformer: $94.99\%$) and graph-based GCAE ($94.79\%$) achieve comparable performance on OSM-O dataset with building footprints without geometric transformations. Conversely, the convolutional VeerCNN shows lower performance of $90.54\%$. When testing on OSM-R with randomly rotated and reflected footprints, PolyMP maintains a robust performance of $85.62\%$ and the baseline methods suffer performance deterioration at various levels (VeerCNN: $90.54\%\rightarrow 78.51\%$, DeepSet: $95.71\%\rightarrow 84.86\%$, SetTransformer: $94.99\%\rightarrow 83.87\%$, GCAE: $94.79\%\rightarrow 80.19\%$).
Performances of models fine-tuned on OSM footprints with original spatial coordinates drop for all models except for PolyMP, which maintains a performance of $81.58\%$ O.A.. This robust performance of PolyMP against the translation of spatial coordinates can be accounted for by the message-passing network of PolyMP where the local geometric features of polygons are computed (i.e., the relative positions of polygons $msg(x_{i},x_{j})=|x_{j}-x_{i}|$). This relative positional features reduce the impact of absolute positioning and effectively normalise the spatial coordinates of polygons while preserving the geometric properties of polygons. VeerCNN outperforms DeepSet ($24.34\%$ O.A) and CGAE ($23.18\%$ O.A) and achieves a test performance of $42.68\%$ O.A. when fine-tuned on the original OSM footprints. We account for VeerCNN’s performance by the built-in translation invariant priors in conventional convolution neural networks. Similarly, SetTransformer shows comparable performance as VeerCNN ($44.72\%$ O.A), which can be attributed to the self-attention features based on the fully connected graph of point sets.
In Fig. 7, we visualise the classification results of PolyMP evaluated on spatial building footprints in OSM-O (Fig. 7 (a)) and OSM-R (Fig. 7 (b)). Predictions indicate that the fine-tuned PolyMP can robustly classify building footprints on a map in varying sizes and geometric shapes based on the geometric invariant features learnt by message-passing networks, taking the graph-based representation of polygons as inputs.
Table 5: Fine-tune test performances on OSM dataset. OSM-O and OSM-R acc. suggest the the classification accuracy of models tested on original and rotated & reflected samples. Overall accuracy is reported for OSM O.A..
@@
VeerCNN 90.54 78.51 84.52 Normalised coordinates $\in(-1,1)$
DeepSet 95.71 84.86 90.56
SetTransformer 94.99 83.87 89.42
GCAE 94.79 80.19 88.02
PolyMP 98.04 85.62 91.92
VeerCNN 42.35 43.01 42.68 Original coordinates
DeepSet 24.12 24.08 24.34
SetTransformer 44.67 44.77 44.72
GCAE 23.44 23.04 23.18
PolyMP 89.86 73.56 81.58
@@
Figure 7: Shape classification of building footprints by PolyMP on the selected region in OSM dataset (downtown areas of Shanghai, China [6]). Model predictions are colored in red.

### v2

@@ 3.2 Message-passing Encoder
where $\{x_{i},x_{j}\}\in A=1$ and $msg(\cdot)$ is the message function that takes the source and target nodes as inputs. From Eq.3, the node feature of $x_{i}$ is updated with the computed message $(x_{i},x_{j})$ and transformed linearly into latent space, producing the latent feature $h_{i}\in\mathcal{H}$. The permutation-invariant message-passing encoder is represented in vector form as:
@@
, where a global readout function (e.g., mean or max pooling) aggregates the latent node-wise features ($\mathcal{H}$) and computes a global graph embedding by applying a non-linear transformation to the pooled features via $\rho_{\theta}$.
Building on geometric message-passing methods for point clouds [36, 11] and polygons [35], we define a specific message-passing function $msg(\cdot)$ for PolyMP as:
$\begin{array}[]{c}msg(x_{i},x_{j})=\alpha*\dfrac{(x_{j}-x_{i})}{std_{\mathcal{N}_{x_{i}}}+\epsilon}+\beta.\end{array}$ (5)
, where $\alpha$ and $\beta$ are learnable parameters that adjust the local neighborhoods to a normal distribution, and $std_{\mathcal{N}_{x_{i}}}$ computes the standard deviation of the local neighborhoods $\mathcal{N}{x_{i}}$ with a small $\epsilon$ to ensure numerical stability. This operation facilitates stable geometric feature learning of graph neighborhoods across diverse geometric structures [39], accommodating the varying properties of polygons (Fig.5). After the message-passing encoder layers, a global pooling layer aggregates the node features to generate a whole-graph embedding, which is then used for downstream polygon shape classification.
@@
Figure 1: Model architecture of PolyMP and PolyMP-DSC. PolyMP consists of a simple message-passing encoder followed by a classification layer (multi-layer perceptron). Yellow arrows indicate the densely self-connected message-passing in PolyMP-DSC.
Following Eq. 3-5, we introduce a simple Message-passing neural network for Polygon geometries, named PolyMP, as shown in Fig. 1. PolyMP processes polygon point coordinates as input node features and the connectivity of polygon edges as input edges. The message-passing encoders of PolyMP compute the “message” for each node based on the graph adjacency and update the node features in each layer.
3.3 Densely Self-Connected Message-Passing
As noted in recent studies of graph message-passing networks [40], the basic message-passing mechanism restricts information flow due to the absence of self-loop information in graph representation learning. To address this limitation and enhance PolyMP’s learning capability, we adopt a densely self-connected (DSC) message-passing mechanism [41, 40, 18], defined as:
$\begin{array}[]{c}\mathcal{G}_{l}=f(\rho_{\theta}(readout(h_{i}|h_{i}\in\mathcal{H}_{l}))+\mathcal{G}_{l-1}+\mathcal{G}_{l-2}+...+\mathcal{G}_{0}),\end{array}$ (6)
where we extend the basic PolyMP model to PolyMP-DSC, as shown in Fig. 1. The PolyMP-DSC aggregates latent graph-wise features $\mathcal{G}_{l}$ at layer $l$ by incorporating latent graph-wise features $\mathcal{G}_{l-1...0}$ from previous message-passing layers. This enables the model to capture hierarchical self-looped graph information, resulting in richer graph-wise representations.
3.4 Implementation
@@
NUFTSpec [8] 2 {288, 128, 64} 48,635 -
DSC-NMP [40] 2 {2, 64, 64} 20,539 Global add.
PolyMP (Ours) 2 {2, 64, 64} 11,837 Global max
PolyMP-DSC (Ours) 2 {2, 64, 64} 16,579 Global max
We evaluate the performance of learning models from previous studies, specifically: the VeerCNN model from Veer et al. [2], which uses 1D sequences converted from polygons as input; the DeepSet model from Zaheer et al. [9] and SetTransformer model from Lee et al. [12], which treat polygon vertices as input point clouds; the GCAE model from Yan et al. [6], which represents polygons as input graphs; and the DSC-NMP model from Fan et al. [40], a baseline graph learning model that processes polygons as input graphs and applies DSC neural message-passing for graph feature learning. We also evaluate the NUFTSpec model from Mai et al. [8], which encodes polygon geometries by converting them into feature vectors in the rasterized spectral domain. These models are compared against the proposed graph message-passing models for polygons, PolyMP and PolyMP-DSC.
The model configurations used in our experiments are summarized in Table 1, and the data representations of polygons and baseline learning models are illustrated in Figure 2. For experimental control, every model trained and tested was constrained to a comparable model depth and latent space of feature embedding. During model training on the Glyph dataset and fine-tuning on the OSM dataset, we compute cross-entropy loss [42] to determine training gradients for fast convergence and good model generalizability. We use Adam optimization [43] with an initial learning rate of 0.01. For better model generalization, we adopt a learning rate decay strategy that reduces the learning rate by a factor of 10 once the model training reaches a performance plateau after 25 epochs. An early-stopping mechanism halts the training once the loss ceases to decrease for 50 epochs, mitigating overfitting. The batch size is set to 64, and the total training epochs are set to 100. Importantly, the models are not designed to individually outperform state-of-the-art models for each architecture, but rather to maintain comparability across architectures and enable the evaluation of the effects of data representations on results.
4 Experiment
4.1 Glyph Dataset
We introduce a synthetic dataset of highly variable geometric shapes to benchmark the classification performance of learning models on vector polygons. This dataset consists of 26 Latin alphabet character glyph geometries, representing semantic classes from A to Z, gathered from an online source [44]. A similar approach has been used previously in computational geometry to construct rich datasets for algorithm testing [45].
The synthetic dataset includes the boundaries (contour lines) of glyphs extracted from 1,413 sans serif and 1,002 serif fonts, resulting in 2D simple polygon geometries compliant with the standards of OpenGeospatialConsortium [38]. Serif and sans serif fonts are the two primary typographic families. Serif glyphs feature decorative strokes that enhance the legibility of body text, while sans serif glyphs have clean, minimal strokes, making them more suitable for headers. Importantly, these minor variations do not affect the overall Gestalt of the shapes, as readers can consistently recognize the symbols. This stability ensures that the labels assigned to the symbols are both stable and robust.
Each polygon geometry is encoded into a fixed-size feature matrix of size $\in\mathbb{R}^{n\times 3}$, where each feature vector $\in\mathbb{R}^{3}$ contains the 2D coordinates of vertices $(x,y)$ as the geometric feature, and a binary feature $(0,1)$ to indicate the position of each 2D coordinate (whether it lies on the outer rings or holes of the polygons). Examples of glyph geometries are shown in Figure 3.
4.2 OSM Dataset
Latin alphabet glyph geometries share geometric similarities with building footprints. To test the generalizability of learning models on spatial objects, we use a building geometry dataset proposed by Yan et al. [6]. The dataset contains 10,000 real-world building footprints extracted from OpenStreetMap (OSM) [23], labeled into 10 categories based on template matching to letters [46]. The building footprints are randomly rotated and reflected from the original canonical samples, as shown in Figure 4. This dataset is currently the most comprehensive real-world collection of vector geometries with purely shape-based labels (as opposed to labels based on building use, which should not be used for shape classification). However, we note that the assignment of labels in this dataset is somewhat subjective and may be less robust compared to the glyph-based dataset. Specifically, shape reflection can pose challenges for label stability, especially for asymmetric letters (e.g., $R$) or mirror image shapes of distinct labels (e.g., $S$ and $Z$).
4.3 Pre-processing
Label-preserving transformations
To assess the invariance of geometric transformations in polygon representations and their effect on learning models, we apply label-preserving geometric transformations to the glyph polygons. We create four transformed training datasets, each consisting of 20%, 40%, 60%, and 80% of glyph polygons that have undergone these transformations. The details of the transformations are listed in Table 2, with the original and transformed data samples visualized in Fig. 5. This results in five training datasets: 0%, 20%, 40%, 60%, and 80% transformations, which will be used for the experiment. The test set consists of data randomly sampled from these five datasets.
Transformation Operation
Polygon Simplifications
To assess the invariance of polygon representations to trivial vertices, we apply the Douglas–Peucker algorithm [47] with a tolerance of 1.0 to the original glyph polygons (typically sized around 50 $\times$ 50 units). This simplification process generates polygon samples with fewer vertices, particularly by excluding co-linear or nearly co-linear vertices. Importantly, the simplification is applied before normalizing the polygons to the range of (-1, 1) and before any augmentation, ensuring that the simplified geometries preserve the topology and essential shape of the original glyph (see Fig. 6).
@@ 5.1 Results on Glyph Dataset
NUFTSpec 96.15 84.40 89.24 83.46 88.32
DSC-NMP 94.31 74.31 86.75 78.36 83.46
PolyMP 98.48 91.64 94.69 92.36 94.31
PolyMP-DSC 98.93 93.19 95.95 94.26 95.61
@@
NUFTSpec 96.10 84.56 89.16 83.28 88.27
DSC-NMP 93.06 73.10 85.03 76.99 82.08
PolyMP 97.42 90.55 93.25 91.36 92.97
PolyMP-DSC 98.47 92.78 95.10 93.43 94.77
Table 3 and Table 4 present the model performances on the Glyph dataset and simplified polygons, respectively.
At the $0\%$ transformation ratio in Table 3, we observe that PolyMP achieves $99.68\%$, $39.91\%$, and $82.35\%$ accuracy on the original, rotated, and scaled test samples, respectively, while VeerCNN achieves $59.07\%$ accuracy on sheared samples. In comparison, DeepSet, SetTransformer, and GCAE maintain comparable performance on the original and scaled test samples ($94.62\%$, $65.01\%$; $96.25\%$, $67.91\%$; and $95.16\%$, $64.54\%$, respectively). However, models that learn from set (DeepSet and SetTransformer) and graph inputs (GCAE and DSC-NMP) experience significant performance deterioration on rotated and sheared samples, with accuracy of $23.67\%$, $26.00\%$; $25.23\%$, $27.79\%$; $31.38\%$, $38.84\%$; and $35.71\%$, $43.21\%$, respectively. Overall, PolyMP-DSC, which learns densely connected graph features, achieves the highest classification accuracy at $71.21\%$, significantly outperforming the baseline DSC-NMP model ($62.24\%$ O.A.).
At the 20% transformation ratio, both PolyMP (84.89% (+44.98%), 93.26% (+10.91%), and 88.17% (+34.95%)) and PolyMP-DSC (86.20% (+44.46%), 94.60% (+8.15%), and 90.65% (+34.49%)) show significant improvements in test accuracy on rotated, scaled, and sheared samples, respectively. In terms of overall accuracy, PolyMP achieves 91.30% O.A. (+22.27%), outperforming VeerCNN’s 81.05% O.A. PolyMP-DSC records the highest O.A. at 92.64% (+21.43%). In contrast, the baseline learning models are unable to match the performance of PolyMP and PolyMP-DSC, despite the increase in training set diversity.
At the 40% and 60% transformation ratios, we observe that the accuracy improvements from increasing the proportion of geometrically transformed samples in the training set have plateaued. In this case, PolyMP achieves 93.11% and 93.69% O.A., while PolyMP-DSC leads with 94.18% and 94.91% O.A. Both models maintain their superior performance compared to others. VeerCNN records 83.30% and 84.88% O.A., followed by DeepSet (80.12% and 81.49% O.A.), SetTransformer (80.33% and 81.94% O.A.), GCAE (72.72% and 75.50% O.A.), DSC-NMP (57.97% and 67.73% O.A.), and NUFTSpec (70.74% and 78.71% O.A.), all of which fall behind.
At the 80% transformation ratio, PolyMP and PolyMP-DSC achieve the highest test performances on rotated and sheared samples, with accuracy of 91.64%, 92.36%, and 93.19%, 94.26%, respectively. These empirical results suggest that, compared to sequence encoding of polygons using CNNs and set representations with set-based learning models, graph representations combined with message-passing neural networks (PolyMP and PolyMP-DSC) are more robust to permutations of polygon vertices. This robustness is particularly evident in geometric transformations such as rotation and shearing, where the feed-in order of vertices can vary. Comparing basic PolyMP with the extended PolyMP-DSC, we observe testing performance improvements of up to $2\%$ O.A. on the Glyph dataset. These results highlight the effectiveness of incorporating hierarchical self-looped graph representations via densely self-connected message passing for polygon encoding and shape-based classification.
From a different perspective, as the proportion of original samples progressively decreases in the training set (from 0% to 80% transformation ratios), PolyMP and PolyMP-DSC experience only a slight test performance deterioration, with a decrease from 99.68% to 98.48% (-1.20%) and 99.58% to 98.93% (-0.65%), respectively. In contrast, VeerCNN (-4.13%), DeepSet (-3.96%), SetTransformer (-5.08%), GCAE (-5.58%), NUFTSpec (-2.27%), and DSC-NMP (-2.50%) exhibit more significant performance drops in test accuracy on the original samples.
In Table 4, we present the test performances of models trained and evaluated on simplified polygons from the Glyph dataset. This experimental setting focuses on evaluating the models’ performance on polygons with a reduced number of trivial vertices (Fig. 6) and compares these results with those shown in Table 3.
We observe similar performance trends in the models when evaluated on the simplified polygons. Notably, VeerCNN experiences a significant drop in performance compared to the results in Table 3. Specifically, at the 0% transformation ratio, VeerCNN’s test performance decreases from 69.43% O.A. (Table 3) to 66.56% O.A. (Table 4), representing a 2.87% decline. As data diversity increases, the performance deterioration in VeerCNN becomes more pronounced, with a drop of 5.2%, progressively increasing to 5.56% O.A. at transformation ratios from 20% to 80%.
The performance deterioration of VeerCNN strongly suggests that CNN models are not robust to changes in the number of vertices in polygons. This can be attributed to several factors, including the data encoding method used in VeerCNN, which represents polygons as fixed-length 1D sequences with zero padding. Simplifying polygons by removing trivial vertices alters the length of these sequences, which not only dilutes the geometric information through zero padding but also disrupts the explicit encoding of vertex adjacency within the 1D sequences.
In comparison, we observe only minor performance drops in general set-based and graph-based learning models. This can be attributed to the flexibility of set and graph representations of polygons, which allow these models to accept input polygons of varying lengths. Specifically, graph-based message-passing networks (i.e., PolyMP and PolyMP-DSC) effectively learn local geometric features from neighboring nodes using a message-passing mechanism, followed by permutation-invariant pooling operations (e.g., max-pooling), as shown in Eq. 4. A similar pooling operation is utilized in set-based learning models [9, 12].
@@ 5.2 Results on OSM Dataset
Table 5: Fine-tuned test performance on the OSM dataset. OSM-O and OSM-R indicate the classification accuracy of models tested on original and rotated & reflected samples, respectively. Overall accuracy is reported as OSM O.A.
@@
VeerCNN 86.30 72.44 79.36 Normalised coordinates $\in(-1,1)$
DeepSet 94.79 82.31 88.62
SetTransformer 89.10 68.73 78.90
GCAE 95.23 78.15 86.94
NUFTSpec 88.02 60.66 74.26
DSC-NMP 87.58 66.73 77.14
PolyMP 96.59 81.03 88.58
PolyMP-DSC 95.75 78.63 87.20
VeerCNN 42.23 42.09 42.16 Original coordinates
DeepSet 22.00 21.65 21.84
SetTransformer 35.62 35.30 35.46
GCAE 22.12 22.32 22.22
NUFTSpec 88.02 61.38 74.42
DSC-NMP 31.81 33.23 32.52
PolyMP 50.64 39.86 46.20
PolyMP-DSC 44.19 36.78 40.68
The results in Table 5 show the test accuracy of fine-tuned classifiers on real-world building footprints from the OSM dataset, using feature encoders pre-trained on the Glyph dataset (80% subset). These fine-tuned models were evaluated on building footprints with both normalized and original coordinates.
PolyMP achieves an accuracy of 88.58% for building footprints with normalized coordinates and 46.20% for footprints with original coordinates. PolyMP-DSC demonstrates slightly lower test accuracy, with 87.20% for normalized coordinates and 40.68% for original coordinates, but still outperforms other baseline methods, except for DeepSet. This can be attributed to the ability of both models to compute local geometric features via message-passing networks, reducing the impact of absolute positioning while preserving polygon geometry. DeepSet performs best, with 88.62% O.A. on normalized coordinates and 82.31% O.A. on rotated and reflected buildings, making it the top performer in some cases.
The NUFTSpec model, which transforms polygon geometries into consistent spectral feature vectors via Non-Uniform Fourier Transforms, demonstrates stable performance at approximately 74% O.A. across both normalized and original coordinates. This stability stems from a series of affine transformations (e.g., scaling and translation) applied to input polygons during feature transformation, ensuring that each polygon is projected into a unit space and positioned within the same relative coordinate system, as discussed in [8].
In Fig. 7, classification results of PolyMP on building footprints from the two OSM subsets (OSM-O and OSM-R) are visualized. The results show PolyMP’s robustness in classifying building footprints of varying sizes and geometric shapes, leveraging the geometric invariant features learned through message-passing networks.
These findings underline the effectiveness of graph message-passing models, particularly PolyMP and PolyMP-DSC, in handling both normalized and original coordinate systems and achieving strong performance across various data transformations.
@@
Figure 7: Shape classification of building footprints using PolyMP on a selected region of the OSM dataset [6]. Model predictions are highlighted in red.